OpenStudy (anonymous):

Let . f(x)=2x+3 The rate of change of f(x) over any interval equals 3. A. True B. False

OpenStudy (anonymous):

what is the change from f(1) to f(2) ? then answer what is the change from f(2) to f(3)?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

am i correct?

OpenStudy (anonymous):

no answer this: f(1) = f(2) = f(3) = f(4) =

OpenStudy (anonymous):

see how those values are changing for each 1 unit change in x.

OpenStudy (anonymous):

yes! :) so what does that mean? LOL

OpenStudy (anonymous):

the change from f(1) to f(2) is your answer.

OpenStudy (anonymous):

WAIT! SO THE ANSWER IS FALSE BECAUSE THEY CHANGE THERE NOT GOING TO ALWAYS BE 3! right?

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

correct... the change is 2

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

good job man :)